# Geological and Well prior assisted full waveform inversion using conditional diffusion models

### v1

@@ 4 Field data application
In this section, we will apply our method to field data to further validate the effectiveness of the geological and well-prior conditioned diffusion regularized FWI. We specifically test the proposed method on 2D marine streamer data acquired from the North-Western Australian Continental Shelf by CGG. To reduce the computational cost of FWI, we select 116 shots with a lateral interval of 98.75 $m$ from the original 1824 shots with an interval of 18.75 $m$. The receiver gather includes 648 receivers with a spacing of 12.5 $m$, yielding a minimum and maximum offset of 16.9 $m$ and 8256 $m$, respectively. The data were recorded for a maximum duration of about 7 seconds with a 2-$ms$ sampling interval. The velocity model covers an area of 12.5 $km$ in length and 3.7 $km$ in depth, with a grid interval of 12.5 $m$ in both $x$ and $z$ directions. Similar to our previous preprocessing (Wang et al., 2023), we first applied a low pass filter on the field data to filter out the component above 8.0 Hz. The source wavelet is inverted from the near-offset early arrival seismograms with a constant water velocity of 1500 m/s. Then, the following sequences of processing steps, including a brute-stack time-domain velocity analysis (Kalita and Alkhalifah, 2019) followed by an extended full waveform inversion with matching filter (Li and Alkhalifah, 2021), were used to obtain a good initial model (shown in Figure 15a). To highlight the influence of priors with different modalities on the inverted results using our method, we only test the single prior conditioned diffusion FWI in this section.
@@
Figure 12: The inverted results using the patch-based class conditioned diffusion regularized FWI. The result with a class condition of "FlatFaultB" (a), that with "CurveFaultA" (b), and that with "Customized" (c). The arrows point to areas of differences in which (a) and (b) show more fault features and artifacts.
@@ 4.1 Geological prior assisted FWI
We first test the geological prior assisted FWI on the field data. Since we do not have an exact geological description of the field subsurface, we created the datasets of velocity models with relatively flat or mild curvature features inside the model and with a resolution of $152\times 152$, marked as a class "Customized". Then, we combine this class with two classes from the OpenFWI dataset: "CurveFaultA" and "FlatFaultB" (for which we upsample the resolution to $152\times 152$) to have a total of three classes in our training data. We train a conditional diffusion model using these three classes with the same training configuration mentioned above. Because the training of pixel space diffusion model on high-resolution images, like the model size considered for this FWI, is hard and costly, instead, we train the diffusion model using a resolution of 152$\times$152 and then use patch-based strategies during the FWI (Wang et al., 2023). The sampled results using the trained diffusion model conditioned on three classes are shown in Figure 11.
Here, we use 11 reverse diffusion steps and do 8 iterations of FWI within each diffusion step. The step length is 0.002 for the FWI update. As the size of the velocity model is 296$\times$1000 grid points, directly inputting the velocity model into the diffusion model originally trained on the resolution of 152$\times$152 will result in artifacts. Instead, we use a patch-based strategy to extract six horizontal patches of size 296$\times$296 and then merge them back to the original velocity model size. The results using different geological priors are shown in Figure 12. Although the initial velocity model has already provided sufficient information on the background to FWI, different geological priors will result in different inverted results. The class label "FlatFaultB" promotes more flatness and lateral discontinuity (Figure 12a) and turns the curved areas into faults. Using the class label "CurveVelA" as a condition, more curvatures (Figure 12b) appear in the inverted result, but the artifacts in the horizon direction are hard to suppress. When using the class label "Customized", the higher wavenumber components of the subsurface model are reconstructed with better horizontal continuity, as shown in Figure 12c. With this better geological prior, the conditional diffusion regularized FWI achieves the best result.
@@ 4.2 Well conditioned diffusion regularized FWI
After we have tested the trained diffusion model to generate velocities consistent with the trained distribution and close to the provided well information, we use this model in FWI. Similar to the geological prior conditioned diffusion regularized FWI test, we split the subsurface velocity model into 6 patches and give only the patches that cover the well log location with the condition scale $\lambda$ equal to 19 (denoting conditional generation), while others with -1 (denoting the unconditional generation). For unconditional diffusion FWI, the condition scale is equal to -1 for all patches. As shown in Figure 15, since the initial model is the result of an extended FWI using a matching filter, the conventional FWI is generally able to converge, providing mainly (or only) high-resolution information (Figure 15b). In contrast, the diffusion-regularized FWI and the well-conditioned diffusion-regularized FWI enhance the lateral continuity of the inverted model and remove the illumination artifacts at the edge of the velocity model. Particularly for the result using the well-conditioned FWI, it recovers the higher wavenumber components of the velocity model. We extract the profiles at the well location at a lateral position of 10.5 $km$ for comparison, which is shown in Figure 16. In general, the profile of the inverted result using the proposed method is closer to the well log than the one obtained using the conventional FWI as well as the diffusion regularized FWI. Especially at the depth range from 1.5 to 2.8 $km$, the result using our method has obvious improvements. Similar to the results shown for the synthetic tests, the well-log guidance can somehow adjust the velocity value closer to the well log and match the fluctuations in the velocity profile. However, we do observe some parts at a depth of 3.0 km, compared to other methods, show lower velocity values even though the velocity fluctuations matched much better. This reason is that our training dataset has a low-velocity value in this region, resulting in velocity bias (preference) to lower values around this depth. This demonstrates the ability of the diffusion model to store the distribution of the training velocity models, and such distribution will guide the inversion as it represents the prior distribution. In addition, we can observe the effectiveness of the well-prior conditioned regularized FWI, and it generally performs better than other methods.
@@
Figure 15: Starting from the initial velocity model (a), we obtain the inversion results from (b) conventional FWI, (c) diffusion FWI where the conditional factor for well log is -1 (no well log information), and (d) well-conditioned diffusion FWI where the conditional factor for well log is 19. Compared with the inversion result in (b) and (c), the well-conditioned diffusion FWI can admit higher wavenumber components in (d), especially in the areas denoted by the arrows.
@@ 5 Discussion
We used simple geological classes and well velocities to condition our diffusion-based generator prior. The results demonstrated the power of such conditioning in guiding the FWI results. There are no limits to including various information in the guidance, granted they are included in the training. Using more realistic models for training (industry-level models, not academia), we can include more high-resolution geological details like thin layers, salt bodies, and facies and even specify their potential locations by incorporating localized classes. A form of such details can be accomplished by including facies maps along with the velocity models in training the diffusion models. For more structural guidance, we can include a migrated image obtained using the initial model as a condition. In summary, the potential for including prior information and interpreter input into FWI is limitless using diffusion models.
@@
Regarding the training velocity models, we generate 20000 samples with a resolution of $500\times 128$ guided by the well logs and structural prior (Ovcharenko et al., 2022) (remove the water layer), and some samples of the training dataset are shown in Figure 17. We use the same configuration as the OpenFWI test to train the conditional diffusion model, except we increase the size of the embedding layer for the well log embedding as we hope to incorporate high-resolution information in the wells. Figure 18 shows 16 samples obtained by conditional diffusion models given different well-log profiles. The matching between the well-log profiles and profiles extracted from the predicted velocity models is not perfect. This drawback is coming from the challenge of the diffusion model to handle high-resolution images (Rombach et al., 2022). Compared to the training samples (Figure 17), although the samples (Figrue 18a) fit the structure information and velocity changes, from the comparison of the well profiles, the prediction only matches the trend of the well log.
@@
Then, we further apply the proposed method to invert for the velocity, shown in Figure 19. We observed the influence of the well condition on the inverted velocity even though the region is far away from the well log location, especially at the left part of the velocity model, whose wavenumber is higher than the conventional one (Figure 15b) and has velocity consistency with the right parts. We also extract the profiles at the well location of 10.5 $km$ for comparison, shown in Figure 20. Although the proposed method with a high-resolution diffusion model does not perform better than the one using a patch-based strategy, it is closer to the well log than the one obtained using the conventional FWI and matches the trend of the well log. To further enhance the proposed method with a diffusion model, we argue that utilizing a latent diffusion model (Rombach et al., 2022) is beneficial, which we will attempt to do in the future.

### v2

@@ 4 Field data application
In this section, we applied our method to field data to further validate the effectiveness of the geological and well-prior conditioned diffusion regularized FWI. We specifically test the proposed method on 2D marine streamer data acquired from the North-Western Australian Continental Shelf by CGG. To reduce the computational cost of FWI, we select 116 shots with a lateral interval of 98.75 $m$ from the original 1824 shots with an interval of 18.75 $m$. The receiver gather includes 648 receivers with a spacing of 12.5 $m$, yielding a minimum and maximum offset of 16.9 $m$ and 8256 $m$, respectively. The data were recorded for a maximum duration of about 7 seconds with a 2-$ms$ sampling interval. The velocity model covers an area of 12.5 $km$ in length and 3.7 $km$ in depth, with a grid interval of 12.5 $m$ in both $x$ and $z$ directions. Similar to our previous preprocessing (Wang et al., 2023), we first applied a low-pass filter on the field data to filter out the component above 8.0 Hz. The source wavelet is inverted from the near-offset early arrival seismograms with a constant water velocity of 1500 m/s. Then, the following sequences of processing steps, including a brute-stack time-domain velocity analysis (Kalita and Alkhalifah, 2019) followed by an extended full waveform inversion with matching filter (Li and Alkhalifah, 2021), were used to obtain a good initial model (shown in Figure 15a). To highlight the influence of priors with different modalities on the inverted results using our method, we only test the single prior conditioned diffusion FWI in this section.
@@
Figure 12: The inverted results using the patch-based class conditioned diffusion regularized FWI. The result with a class condition of "FlatFaultB" (a), that with "CurveFaultA" (b), and that with "Customized" (c). The arrows point to areas of differences in which (a) and (b) show more fault features and artifacts. These results demonstrate that conditional diffusion FWI improves inversion accuracy and that incorporating suitable prior information as conditioning leads to further enhancements.
@@ 4.1 Geological prior assisted FWI
We first test the geological prior assisted FWI on the field data. Since we do not have an exact geological description of the field subsurface, we created the datasets of velocity models with relatively flat or mild curvature features inside the model and with a resolution of $152\times 152$, marked as a class "Customized". Then, we combine this class with two classes from the OpenFWI dataset: "CurveFaultA" and "FlatFaultB" (for which we upsample the resolution to $152\times 152$) to have a total of three classes in our training data. The U-Net model has demonstrated strong capacity across different OpenFWI datasets, which encompass a variety of geological complexities such as curvatures and fault structures. For the field data case, we adopt a patch-based approach, and we found that the current U-Net architecture has sufficient capacity to effectively handle localized structures within each patch. Hence, we train a conditional diffusion model using these three classes with the same training configuration mentioned in the synthetic tests. Because the training of pixel space diffusion model on high-resolution images, like the model size considered for this FWI, is hard and costly, instead, we train the diffusion model using a resolution of 152$\times$152 and then use patch-based strategies during the FWI (Wang et al., 2023). The sampled results using the trained diffusion model conditioned on three classes are shown in Figure 11.
Here, we use 11 reverse diffusion steps and do 8 iterations of FWI within each diffusion step. The step length is 0.002 for the FWI update. As the size of the velocity model is 296$\times$1000 grid points, directly inputting the velocity model into the diffusion model originally trained on the resolution of 152$\times$152 will result in artifacts. Instead, we use a patch-based strategy to extract six horizontal patches of size 296$\times$296 and then merge them back to the original velocity model size. The results using different geological priors are shown in Figure 12. Although the initial velocity model has already provided sufficient information on the background to FWI, different geological priors will result in different inverted results. The class label "FlatFaultB" promotes more flatness and lateral discontinuity (Figure 12a) and turns the curved areas into faults. Using the class label "CurveVelA" as a condition, more curvatures (Figure 12b) appear in the inverted result, but the artifacts in the horizon direction are hard to suppress. When using the class label "Customized", the higher wavenumber components of the subsurface model are reconstructed with better horizontal continuity, as shown in Figure 12c. With this better geological prior, the conditional diffusion regularized FWI achieves the best result. Further waveform misfit analysis (Figure 17) will demonstrate this observation as well.
@@ 4.2 Well conditioned diffusion regularized FWI
After we have tested the trained diffusion model to generate velocities consistent with the trained distribution and close to the provided well information, we use this model in FWI. Similar to the geological prior conditioned diffusion regularized FWI test, we split the subsurface velocity model into 6 patches and give only the patches that cover the well log location with the condition scale $\lambda$ equal to 19 (denoting conditional generation), while others with -1 (denoting the unconditional generation). For unconditional diffusion FWI, the condition scale is equal to -1 for all patches. As shown in Figure 15, since the initial model is the result of an extended FWI using a matching filter, the conventional FWI is generally able to converge, providing mainly (or only) high-resolution information (Figure 15b). In contrast, the diffusion-regularized FWI and the well-conditioned diffusion-regularized FWI enhance the lateral continuity of the inverted model and remove the illumination artifacts at the edge of the velocity model. Particularly for the result using the well-conditioned FWI, it recovers the higher wavenumber components of the velocity model. We extract the profiles at the well location at a lateral position of 10.5 $km$ for comparison, which is shown in Figure 16. In general, the profile of the inverted result using the proposed method is closer to the well log than the one obtained using the conventional FWI, as well as the diffusion regularized FWI. Especially at the depth range from 1.5 to 2.8 $km$, the result using our method has obvious improvements. Similar to the results shown for the synthetic tests, the well-log guidance can somehow adjust the velocity value closer to the well log and match the fluctuations in the velocity profile. However, we do observe that in some deeper regions, at a depth of 3.0 km, the inverted velocity from the conditional diffusion FWI appears biased toward lower values, despite showing improved alignment in terms of velocity fluctuations. This discrepancy arises because the training dataset contains relatively low-velocity values in this depth range, which introduces a bias in the learned prior distribution. Since diffusion models implicitly capture and sample from the data distribution they are trained on, this prior preference influences the inversion outcome, even when well-log information is provided as a condition. This behavior highlights a key property of the diffusion model to store the distribution of the training velocity models, and such a distribution will guide the inversion as it represents the prior distribution. In addition, we can observe the effectiveness of the well-prior conditioned regularized FWI, and it generally performs better than other methods. However, it also reveals a limitation: when the training distribution is not sufficiently diverse or representative (e.g., lacking higher velocities at certain depths), the model may be biased toward the dominant patterns in the training set.
@@
Figure 15: Starting from the initial velocity model (a), we obtain the inversion results from (b) conventional FWI, (c) diffusion FWI where the conditional factor for well log is -1 (no well log information), and (d) well-conditioned diffusion FWI where the conditional factor for well log is 19. Compared with the inversion result in (b) and (c), the well-conditioned diffusion FWI can admit higher wavenumber components in (d), especially in the areas denoted by the arrows. The well-log guidance can somehow adjust the velocity value closer to the well log and match the fluctuations in the velocity profile.
@@
4.3 A Misfit Curve Analysis for Field Data
To evaluate the effectiveness of different inversion strategies on real seismic data, we compare the misfit convergence curves obtained from a series of field data FWI experiments (see Figure 17). The conventional FWI method exhibits slow but stable convergence, which is expected due to the high nonlinearity and lack of prior constraints. In contrast, the unconditional diffusion FWI accelerates convergence in early iterations by leveraging learned velocity priors from carefully constructed training set distributions, although it lacks control over geological structure.
The conditional diffusion FWI (CD-FWI) variants, conditioned on different geological class labels (e.g., CurveFaultA, FlatFaultB, and a customized class), demonstrate consistently improved convergence rates and lower final misfit values. This highlights the importance of incorporating geological context during the inversion process. Among them, CD-FWI with the customized class shows the best convergence performance, indicating that tailored conditioning aligned with the target data characteristics can further enhance inversion quality.
Moreover, the CD-FWI method incorporating well-log information achieves the lowest overall misfit at the earlier stages (within the first 50 iterations). This suggests that the integration of multi-modal prior constraints, particularly point-wise velocity constraints from well logs, provides significant benefits in stabilizing the inversion and guiding the optimization toward geologically plausible solutions. In the later iterations, however, the waveform misfit does not reach the lowest values due to anisotropy in the region, leading to an inherent mismatch between the well-log velocity and the inverted FWI velocity. If incorporating the anisotropic FWI or with better well-log data, it could be further improved.
Figure 17: Misfit curves from field data inversions using different FWI strategies. The comparison includes conventional FWI, diffusion-based FWI, conditional diffusion FWI (CD-FWI) conditioned on different geological classes, and CD-FWI with well-log constraints. The vertical axis is plotted on a logarithmic scale to highlight convergence behavior across methods.
@@ 5 Discussion
We used simple geological classes and well velocities to condition our diffusion-based generator prior. The results demonstrated the power of such conditioning in guiding the FWI results. There are no limits to including various information in the guidance, granted they are included in the training. Using more realistic models for training (industry-level models, not academia), we can include more high-resolution geological details like thin layers, salt bodies, and facies, and even specify their potential locations by incorporating localized classes. A form of such details can be accomplished by including facies maps along with the velocity models in training the diffusion models. For more structural guidance, we can include a migrated image obtained using the initial model as a condition. In summary, the potential for including prior information and interpreter input into FWI is limitless using diffusion models.
Regarding the additional cost, the proposed approach introduces minimal computational overhead during inference due to the additional diffusion regularized step. Specifically, each diffusion step takes approximately 0.036 seconds on an NVIDIA A100 GPU, which is negligible compared to the runtime of one iteration of conventional FWI. Although pretraining the diffusion model takes several hours, depending on the size of the training dataset and the number of iterations, it is performed entirely offline and only once.. As a result, the inference-time computational cost of our approach remains efficient and decoupled from the training stage. This separation enables practical deployment of the method without impacting the runtime efficiency of the inversion process.
The proposed approach establishes a practical workflow: one can pre-train a diffusion model on a variety of generalized subsurface structures and subsequently apply it to field cases. The preceding field applications have already demonstrated the generalizability of the method, particularly in the context of class-conditioned diffusion FWI. Here, we further assess the generalizability of the well-conditioned diffusion FWI. Specifically, we apply the pre-trained diffusion model (trained on the OpenFWI dataset) to a portion of the Overthrust model (Figure 18a), which differs significantly from the OpenFWI datasets in that it is not piecewise-constant and exhibits distinct geological features. The observation setup follows that of the CurveFaultB test, with the number of sources and receivers adjusted to 8 and 32, respectively. The initial model is derived by applying a Gaussian smoothing ($\sigma=20$) to the ground truth velocity. Figure 18 presents the inverted results from both conventional FWI and the proposed method. Despite being applied to an out-of-distribution scenario, the proposed approach exhibits strong performance, particularly in capturing deep structural details and accurate velocity magnitudes. This further highlights the robustness and generalizability of the proposed method.
Figure 18: a)The true velocity model and inversion results from b) conventional FWI (whose PSNR, SSIM, and relative L2 error are 28.45dB, 0.78, and 0.05, respectively), c) conditional diffusion FWI with a well located at 0.25 km (whose PSNR, SSIM, and relative L2 error are 28.51dB, 0.80, and 0.04, respectively). The proposed conditional diffusion FWI generalizes well in such an out-of-distribution case.
@@
Regarding the training velocity models, we generate 20000 samples with a resolution of $500\times 128$ guided by the well logs and structural prior (Ovcharenko et al., 2022) (remove the water layer), and some samples of the training dataset are shown in Figure 19. We use the same configuration as the OpenFWI test to train the conditional diffusion model, except we increase the size of the embedding layer for the well log embedding as we hope to incorporate high-resolution information in the wells. Figure 20 shows 16 samples obtained by conditional diffusion models given different well-log profiles. The matching between the well-log profiles and profiles extracted from the predicted velocity models is not perfect. This drawback is coming from the challenge of the diffusion model to handle high-resolution images (Rombach et al., 2022). Compared to the training samples (Figure 19), although the samples (Figrue 20a) fit the structure information and velocity changes, from the comparison of the well profiles, the prediction only matches the trend of the well log.
@@
Then, we further apply the proposed method to invert for the velocity, shown in Figure 21. We observed the influence of the well condition on the inverted velocity even though the region is far away from the well log location, especially at the left part of the velocity model, whose wavenumber is higher than the conventional one (Figure 15b) and has velocity consistency with the right parts. We also extract the profiles at the well location of 10.5 $km$ for comparison, shown in Figure 22. Although the proposed method with a high-resolution diffusion model does not perform better than the one using a patch-based strategy, it is closer to the well log than the one obtained using the conventional FWI and matches the trend of the well log. To further enhance the proposed method with a diffusion model, we argue that utilizing a latent diffusion model (Rombach et al., 2022) is beneficial. It can reduce computational cost while preserving high-quality generative capabilities. More importantly, it offers a structured latent space where disentanglement of features (e.g., geological variations) may be easier to achieve. We leave this extension for future investigation.